OpenStudy (anonymous):

Which temperature conversion is correct? °C = K - 273 K = 273 x °C K = °C - 273 °C = K + 273

OpenStudy (anonymous):

d?

OpenStudy (asib1214):

When you convert °C to Kelvins, you add 273K to the °Clike this T= 25°C + 273K = 298K and when you convert Kelvins to °C, you subtract 273K like this T= 30K - 273°C = -243°C

OpenStudy (anonymous):

so it is D?

OpenStudy (asib1214):

No D is not right,

OpenStudy (anonymous):

ok my other choice is A???

OpenStudy (asib1214):

A is right, because whe nyou want your answer in Kelvins and you have °C, you subtract 273

OpenStudy (asib1214):

Like this ... 25°C + 273K = K 30K - 273°C = C